One pipe can fill a pool 1.25 times faster than a second pipe. When both pipes are opened, they fill the pool in five hours. How long would it take to fill the pool if only the slower pipe is used?
1
Expert's answer
2020-11-05T16:36:40-0500

5(x+1.25x)=a; where x is speed(litres per hour) of slower pipe and a is a volume of pool(liters)

x+1.25x=a/5

2.25x=a/5

x=a/11.25

11.25x=a; what means that slower pipe need 11.25 hours (11 hours 15 minutes) to fill the pool

Answer: 11 hours 15 minutes